Selected balance sheet amounts for Reed Corporation for 2017 follow:

Current Assets: $9,450

Long Term Assets: ?

Total Assets: ?

Current Liabilities: $14,520

Long-term Liabilities: $12,340

Total Liabilities: ?

Equity: $15,990

What is the balance sheet amount for Long-term Assets?

Question 5 options:

$26,860

$33,310

$29,500

$42,850
Answer #1

Total assets = Total liabilities + Equity

Total liabilities = Current liabilities + Long term liabilities

Total liabilities = $14,520 + 12,340 = $26,860

Total assets = $26,860 + 15,990 = $42,850

Long term assets = Total assets - Current assets

Long term assets = $42,850 - 9,450 = $33,400